OpenStudy (anonymous):

ok i need help subtracting x+2 over x squared +4x+3 minus x+1 over x+3

OpenStudy (zzr0ck3r):

\[\frac{x+2}{x^2+4x+3}-\frac{x+1}{x+3}\]?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

factor the bottom of the fraction on the left and tell me what you have

OpenStudy (zzr0ck3r):

factor x^2+4x+3

OpenStudy (anonymous):

(x-4((x+1)

OpenStudy (zzr0ck3r):

x^2+4x+3=0?

OpenStudy (zzr0ck3r):

can you solve this?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

well you need to go back and learn how to solve quadratic equations \[\frac{x+2}{(x+3)(x+1)}-\frac{x+1}{x+3}\] so we need to find a common denom and that will be (x+3)(x+1) so we need to multiply the rhs by (x+1)(x+1) \[\frac{x+2}{(x+3)(x+1)}-\frac{(x+1)(x+1)}{(x+3)(x+1)}\]= \[\frac{x+2}{(x+3)(x+1)}-\frac{x^2+2x+1}{(x+3)(x+1)}\]= \[\frac{x+2-x^2-2x-1}{(x+3)(x+1)}\]= \[\frac{-x^2-x+1}{(x+3)(x+1)}\]